What is the concentration, in g/mL, of a solution made by dissolving 75.0 g of phosphoric acid in 50.0 mL of solution?
1
Expert's answer
2020-06-25T07:50:56-0400

The concentration of the solution can be calculated using the following formula:

c = m / V

where c - concentration, m - mass, V - volume of the solution

From here:

c(phosphoric acid) = m(phosphoric acid) / V(phosphoric acid) = 75.0 g / 50.0 ml = 1.50 g/ml


Answer: 1.50 g/mol